OpenStudy (anonymous):

If f(x) = −4x − 1, find f(−2). −9 −7 −10 7

OpenStudy (lgbasallote):

-4(-2) - 1`

OpenStudy (anonymous):

-4(-2)-1 8-1 7

OpenStudy (anonymous):

plug it in again. -4(-2)-1=8-1=7. Your answer is 7 :)

OpenStudy (anonymous):

just remember you always plug in the value :) hope that helps!